OpenStudy (anonymous):

Looking for some help please. Use the method of substitution to solve the system of linear equations. 9x+3y=3/2 y=2x-2 The solution is ? (Type an ordered pairs)

OpenStudy (anonymous):

sub.s the expression for y in the 2nd equation into the first equation and solve for x

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

Hi, yes here I am.

OpenStudy (anonymous):

I am new to algebra so I am looking for help in breaking it down to understand the equation please.

OpenStudy (anonymous):

See how there's a y in the top equation? and how the bottom equation is solved for y? Take that expression for y (2x-2) and put it into the top equation where the y is... then use algebra to simplify and solve...

OpenStudy (anonymous):

Does that part make sense?

OpenStudy (anonymous):

9x+3(2x-2)=3/2?

OpenStudy (anonymous):

yep:)

OpenStudy (anonymous):

can you help me with the next step?

OpenStudy (anonymous):

distribute the 3 to (2x-2)

OpenStudy (anonymous):

add three or multiply?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

9x+3(2x-2)=3/2? 9x+6x-6=3/2 am I right so far?

OpenStudy (anonymous):

looking good:)

OpenStudy (anonymous):

combine terms... the terms with 'x' in them are like terms.... so add them..

OpenStudy (anonymous):

9x+3(2x-2)=3/2? 9x+6x-6=3/2 15x-6=3/2

OpenStudy (anonymous):

the terms with no 'x' in them are like terms so... combine them, which means that that '-6' is going to have to be moved to the other side of the equation to be combined with the '3/2'

OpenStudy (anonymous):

great so far... so moving the -6 what do you need to do to get rid of it on the left side?

OpenStudy (anonymous):

i'm forgetting what I need to do to get rid of the l eft side...

OpenStudy (anonymous):

add 6 to both sides eh? :)

OpenStudy (anonymous):

15x-6+6=3/2+6?

OpenStudy (anonymous):

yep:)

OpenStudy (anonymous):

almost done

OpenStudy (anonymous):

15x=3/2+6 is this right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

should write the 3/2 +6 as a single term, for your own ease...

OpenStudy (anonymous):

3/2 + 12/2 = ?

OpenStudy (anonymous):

15/2

OpenStudy (anonymous):

good. last step: divide both sides by the coefficient on x ie divide both sides by 15...

OpenStudy (anonymous):

and what do you get for x ?

OpenStudy (anonymous):

1/2?

OpenStudy (anonymous):

good:) now go back and use that result to find y...

OpenStudy (anonymous):

easiest to use it here: y=2x-2 put in '1/2' for 'x'

OpenStudy (anonymous):

got it? any questions?

OpenStudy (anonymous):

y=2x(1/2)-2?

OpenStudy (anonymous):

this seems easy but i get frustrated and lost LOL i think i am making it harder than it is.

OpenStudy (anonymous):

uhm use * to signify multiplication or you're going to get it confused with 'x' at some point...

OpenStudy (anonymous):

yeah it gets pretty easy eventually:) bit of practice is all...

OpenStudy (anonymous):

all good?

OpenStudy (anonymous):

y=2*(1/2)-2?

OpenStudy (anonymous):

did I set that up right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

y=2*(1/2)-2? y=1-2

OpenStudy (anonymous):

yes one more little step...

OpenStudy (anonymous):

y=1

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

negative. ok.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so how would "I get the ordered pair. one is -1

OpenStudy (anonymous):

"ordered pair" means you write your answer as 'open parenthesis' 'x value' 'comma' y value' 'close parenthesis' (x,y)

OpenStudy (anonymous):

(1/2,-1)

OpenStudy (anonymous):

put the results we just got in for 'x' and 'y' though

OpenStudy (anonymous):

yep good

OpenStudy (anonymous):

wow, i can't believe I got it. LOL NEVER would have without you. thank you!

OpenStudy (anonymous):

Sure!